OpenStudy (anonymous):

A pound of raisins costs $2.00, and a pound of almonds costs $5.00. Six pounds of a trail mix contains 2 more pounds of raisins than almonds. What is the cost of the 6 pounds of trail mix?

OpenStudy (mathstudent55):

r = weight of almonds. a = weight of almonds a + r = 6 a + 2 = r a + a + 2 = 6 2a = 4 a = 2 r = 4 That means there are 4 lb of raisins and 2 lb of almonds in 6 lb of trail mix. Now use the price per pound of each to find the total cost.

OpenStudy (anonymous):

Thanx!

OpenStudy (anonymous):

I was going to say something very similar to what you said, math but it would look like I was copying you. Gj.

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

@ml1042903 : Got it ?:)

OpenStudy (anonymous):

Yes!